# Neural Implicit Flow Fields for Spatio-Temporal Motion Mapping

### v1

@@ 4.6 Ablation Study
We perform an ablation study on alternative methods for temporal encoding. In our method, we use a SIREN network to process the temporal input. For comparison, we evaluate two alternative mappings of time $t$ into a temporal feature vector $\mathbf{f}_{t}(t)$:
@@
Method $\downarrow NLL$ reduction (vs Ours) 95% CI
@@ 5 Conclusions
Through experiments on the large-scale ATC dataset, we demonstrated that NeMo-map achieves subsantially higher accuracy (lower negative log-likelihood) than existing MoD baselines, while reducing map building time. Qualitative results further show that the learned flow fields capture multimodality, temporal variations, and environment topology without requiring explicit maps. Ablation studies confirmed the advantage of using SIREN-based temporal encoding over discrete or Fourier alternatives.
@@
In summary, the results highlight continuous MoDs as a practical and scalable tool for modeling human motion dynamics. By combining accuracy, efficiency, and flexibility, the representation offers a powerful prior for downstream tasks such as socially aware navigation, long-term motion prediction, and localization in dynamic environments. In future work, we plan to extend this formulation with online update mechanisms to adapt continuously to evolving crowd behaviors, further bridging the gap toward long-term real-world deployment.

### v2

@@ 4.5 Ablation Study
In this section, we analyze alternative methods for temporal encoding and further include hyperparameter ablations on the SWGMM component number $J$ and the spatial feature grid resolution $H\times W$ in Appendix E.
In our method, we use a SIREN network to process the temporal input. For comparison, we evaluate two alternative mappings of time $t$ into a temporal feature vector $\mathbf{f}_{t}(t)$:
@@
Method NLL $\downarrow$ NLL reduction (vs Ours) 95% CI
@@ 5 Conclusions
Through experiments on the large-scale ATC dataset and the outdoor ETH/UCY dataset, we demonstrated that NeMo-map achieves substantially higher accuracy (lower negative log-likelihood) than existing MoD baselines, while offering practical map building times suitable for large-scale datasets. We also demonstrate that the learned flow fields capture multimodality, temporal variations, and environment topology without requiring explicit maps, and achieve higher performance in the trajectory prediction downstream task. Ablation studies confirmed the advantage of using SIREN-based temporal encoding over discrete or Fourier alternatives.
@@
In summary, the results highlight continuous MoDs as a practical and scalable tool for modeling human motion dynamics. By combining accuracy, efficiency, and flexibility, the representation offers a powerful prior for downstream tasks such as socially aware navigation, long-term motion prediction, and localization in dynamic environments. Despite its advantages, the proposed method has limitations in handling sharp spatial or temporal discontinuities, such as temporary barriers or sudden event-driven changes. Adapting to such changes requires additional observations, which is a limitation shared by existing MoD approaches. In future work, we plan to extend this formulation with online update mechanisms to adapt continuously to evolving crowd behaviors, and to better handle sharp, event-driven discontinuities in motion patterns, further bridging the gap toward long-term real-world deployment.
6 Ethics statement
The dataset used for training are publicly available and fully anonymized, representing persons only as 2D positions without identifiers or visual data. Maps of dynamics further aggregate these trajectories into statistical motion patterns, so no personal information are retained.
7 Reproducibility statement
For reproducibility, we provide the full training and evaluation code together with detailed instructions in the supplementary material. The package includes our main model as well as the variants used in the ablation study. Evaluation results with per-sample NLL values are also attached. In addition, we provide scripts for generating and visualizing maps of dynamics (MoDs).
Appendix A The Use of Large Language Models (LLMs)
Parts of this manuscript were edited with the assistance of LLMs to improve grammar and clarity. All scientific content was written and verified by the authors.
Appendix B Dataset
ATC (Brščić et al., 2013)
[…]
ETH (Pellegrini et al., 2009)/UCY (Lerner et al., 2007)
The ETH/UCY datasets contain pedestrian trajectories captured in outdoor environments such as university campuses and shopping areas. We evaluate using all four scenes: ETH and HOTEL from the ETH dataset, and UNIV and ZARA from the UCY dataset.
Appendix C Implementation Details
The output of the network $\Phi_{\theta}$ parameterizes an SWGMM over speed and orientation. For $J$ mixture components, the network predicts $6J$ raw values per query coordinate. Each component $j$ is defined by: a mixture weight $w_{j}$, obtained by applying a softmax over the raw weights; a mean speed $\mu_{j,s}=\max(0,\,\tilde{\mu}_{j,s})$ and mean orientation $\mu_{j,a}=\tilde{\mu}_{j,a}\bmod 2\pi$; variances $\sigma^{2}_{j,s}=\exp(\mathrm{clamp}(\tilde{v}_{j,s},-10,10))$ and $\sigma^{2}_{j,a}=\exp(\mathrm{clamp}(\tilde{v}_{j,a},-10,10))$; and a correlation coefficient $\rho_{j}=0.99\tanh(\tilde{\rho}_{j})$. Altogether, the network defines a valid SWGMM with parameters as in Eq. 3, where $\bm{\mu}_{j}=\big(\mu_{j,s},\,\mu_{j,a}\big)$ and $\bm{\Sigma}_{j}$ is the covariance matrix with diagonal entries $\sigma^{2}_{j,s},\,\sigma^{2}_{j,a}$ and correlation $\rho_{j}$. In the experiments, $J$ is set to 3 and coordinates are normalized to [1,1][-1,1]. Spatial input is processed by an MLP with hidden sizes [128, 64][128,\,64] and ReLU activations. In the ATC dataset, the environment exhibits strong daily regularity and the timestamps are provided in epoch format. We therefore convert them to time-of-day to capture long-term temporal patterns. In the ETH/UCY dataset where no such periodic structure is present, we instead use the frame as the temporal variable. Temporal input is processed by a two-layer SIREN (sine activations with $\omega_{0}^{(1)}=30$ in the first layer and $\omega_{0}^{(h)}=1$ in the hidden layer). The two streams are fused via FiLM modulation (Perez et al., 2018). The fused representation is passed to a linear head producing $6J$ outputs. Models are trained using the Adam optimizer with learning rate $10^{-3}$ for 100 epochs. An ablation of alternative temporal encodings is provided in Sec. 4.5.
Appendix D Downstream task validation
In the experiments of long-term human motion prediction, for observations, we use $3\text{\,}\mathrm{s}$ of each trajectory and use the remaining (up to the maximum prediction horizon) as the prediction ground truth. We compare our method against MoD-based predictors, Trajectory Unified TRansformer (TUTR, Shi et al. (2023)) and Motion Indeterminacy Diffusion (MID, Gu et al. (2022)).
TUTR unifies social interaction modeling and multimodal trajectory prediction components in a transformer encoder-decoder architecture. It predicts multiple trajectories with corresponding probabilities, and we evaluate using the trajectory with the highest probability. MID formulates trajectory prediction as a reverse process of motion indeterminacy diffusion, which gradually discards the indeterminacy to obtain desired trajectory from ambiguous walkable areas. Both MID and TUTR are trained for 100 epochs on both datasets.
For the evaluation metrics, we use Average and Final Displacement Errors (ADE and FDE). ADE describes the mean $L^{2}$ distance between predicted trajectories and the ground truth. FDE describes the $L^{2}$ distance between the predicted and the ground truth positions at the last prediction time step. For each ground truth trajectory we generate $k$ = 5 prediction trajectories. Based on practical applications for autonomous robots, we evaluate using the most likely predicted trajectory. When probability information for the predicted trajectories is not available, such as in the case of the baseline MID, we report the mean ADE and FDE values over the predicted trajectories.
Appendix E Hyperparameter ablation study
Figure 8: Hyperparameter ablations on the SWGMM component number $J$ and the spatial feature grid resolution $H\times W$.
For hyperparameter ablations, we study the effect of the SWGMM component number $J$ and the spatial feature grid resolution $H\times W$. Increasing $J$ allows the model to represent more complex human motion behaviors. However, a larger number of components also increases the output dimension of the MoD representation and leads to higher computational cost during both training and inference. In the left figure of Fig. 8, performance stabilizes when $J\geq 3$, showing that it is sufficient to capture the dominant multimodal motion patterns in human environments. Therefore, we recommend starting with $J=3$ in practical applications.
For the spatial feature grid resolution $H\times W$, a finer grid enables better representation of local motion variations, but also increases memory consumption and training time. As shown in the right figure of Fig. 8, the performance remains stable in the range from $16\times 16$ to $64\times 64$. We therefore recommend starting in this range to have a good trade-off between performance and efficiency.